OpenStudy (sierra!):

What is the value of the expression 14 + 6 • 2 – 10 ÷ 5? A. 38 B. 6 C. 0 D. 24

OpenStudy (welshfella):

to do this you need to know the order of operations PEDMAS Parenthese Exponantials Divide Multiply Add Subtract

OpenStudy (anonymous):

^^

OpenStudy (anonymous):

@sierra! Where do you think we start?

OpenStudy (sierra!):

10/5?

OpenStudy (welshfella):

so for your question do the multiply and divide first then work from left to right

OpenStudy (welshfella):

yes that s ok work out 10/5 first then do the multiplication

OpenStudy (anonymous):

We use the order of PEMDAS P: Parenthesis E: Exponent M/D: Multiplication or division starting from left to right A/S: Addition or subtraction starting from left to right

OpenStudy (anonymous):

We don't have parenthesis or exponents, so what would we do first? @sierra!

OpenStudy (sierra!):

mutiplication

OpenStudy (anonymous):

Yes, so what do you multiply?

OpenStudy (sierra!):

6*2

OpenStudy (sierra!):

=12

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

Then what do we do next?

OpenStudy (sierra!):

divide

OpenStudy (anonymous):

Yes. :)

OpenStudy (sierra!):

All together I got six is that correct

OpenStudy (anonymous):

No...

OpenStudy (sierra!):

D:

OpenStudy (anonymous):

Ok, after you multiply 6 x 2, what is the equation now?

OpenStudy (sierra!):

\[10\div5\]

OpenStudy (sierra!):

my guess is c

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Just follow what I am telling you.

OpenStudy (anonymous):

What is the equation after you multiply 6 x 2?

OpenStudy (sierra!):

Divide

OpenStudy (anonymous):

I said what is the equation...

OpenStudy (anonymous):

*expression

OpenStudy (anonymous):

14+(6)(2)− 10 5 =24

OpenStudy (sierra!):

\[12\div5\]

OpenStudy (sierra!):

thanks

OpenStudy (anonymous):

np:) im glad i could help

OpenStudy (anonymous):

@LightYagami96 Please do not give direct answers. They are against OpenStudy's Code of Conduct. If you keep giving them, you might get suspended. Please refrain from giving direct answers in the future. Thank you. :) If you do not know the rules, then you can read them here: http://openstudy.com/code-of-conduct

OpenStudy (welshfella):

16 + 6 * 2 - 10/5 = 14 + 12 - 10/5 = 14 + 12 - 2 = 26 -2 = 24

OpenStudy (anonymous):

(Facepalm)

OpenStudy (anonymous):

14+(6)(2)− 10 5 =14+12− 10 5 =26− 10 5 =26−2 =24

OpenStudy (anonymous):

That's still a direct answer... it gives out the answer...

OpenStudy (anonymous):

why dont u worry about you and not me? huh!?!?!

OpenStudy (welshfella):

I understood that as long as we explained how to do the problem then it was Ok to give the answer. not so?

OpenStudy (anonymous):

hey

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

@welshfella No, because the asker can just take the answer without looking at what you have explained. Take a look at this:

OpenStudy (anonymous):

@AnswerMyQuestions i helped, is that good? you guys were confusing her/him

OpenStudy (welshfella):

yes - I see the point. Some users questioners will study the method but others will not. Fair point.

OpenStudy (welshfella):

@LightYagami96 - you would have helped if he/she had studied the method but we don't know that

OpenStudy (anonymous):

what are you talking about, you did the same thing.

OpenStudy (anonymous):

We weren't confusing anyone.

OpenStudy (anonymous):

im leaving stop spamming if u wanna talk pm me

OpenStudy (welshfella):

@AnswerMyQuestions - the code of Conduct is not so specific as amistre's post . I think the former needs to be altered accordingly.

OpenStudy (anonymous):

Yes, it isn't very specific.

OpenStudy (welshfella):

right anyway must go